OpenStudy (ja1):

Below:

OpenStudy (jdoe0001):

above

OpenStudy (ja1):

\[\sqrt{169} - \sqrt{50} - \sqrt{8}\]

OpenStudy (ja1):

Hehe

OpenStudy (ja1):

Find the exact value of the radical expression in simplest form.

OpenStudy (ja1):

That's the question.

OpenStudy (jdoe0001):